Phillip has $2.10 in nickels and dimes. If he has a total of 31 coins, how many of the coins are nickels?
Answer by TimothyLamb(4379)

5n + 10d = 210
n + d = 31
d = 31 - n
---
5n + 10d = 210
5n + 10(31 - n) = 210
5n + 310 - 10n = 210
5n = 100
---
n = 20
d = 11
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php